A machine produces 164 bolts in 20 minutes. At the same rate, how many bolts would be produced in 35 minutes?: A machine produces 164 bolts in 20 minutes. At the same rate, how many bolts would be produced in 35 minutes?
Answer by checkley77(3624)

20/35=164/X
35X=20*164
35X=3280
X=3280/35
X=93.7 BOLTS PRODUCED IN 35 MIN.
PROOF:
20/35=164/93.7
35*93.7=20*164
3280=3280